### v1

# CodeAgent: Enhancing Code Generation with Tool-Integrated Agent Systems for Real-World Repo-level Coding Challenges

Abstract
Large Language Models (LLMs) have shown promise in automated code generation but typically excel only in simpler tasks such as generating standalone code units. Real-world software development, however, often involves complex code repositories (named repo) with complex dependencies and extensive documentation. To fill this gap, our research pivots towards evaluating LLMs in a more realistic setting – real-world repo-level code generation. We introduce CodeAgentBench, a manually curated benchmark for repo-level code generation. This benchmark comprises five high-quality Python projects, encompassing a total of 101 samples. We assess nine leading LLMs on repo-level tasks and observe a decline in their performance. To tackle this, we present CodeAgent, a novel LLM-based agent framework that employs external tools for effective repo-level code generation. CodeAgent integrates five programming tools, enabling interaction with software artifacts for information retrieval, code symbol navigation, and code testing. We implement four agent strategies to optimize these tools’ usage. Our experiments on CodeAgentBench show that CodeAgent enhances LLM performance significantly, with improvements ranging from 18.1% to 250%. Further tests on the HumanEval benchmark confirm CodeAgent’s adaptability and efficacy across various code generation tasks. Notably, CodeAgent outperforms commercial products like Github Copilot, showcasing superior accuracy and efficiency. These results demonstrate CodeAgent’s robust capabilities in code generation, highlighting its potential for real-world repo-level coding challenges.  [^1] Code, data, and more case studies are open-sourced on Github: https://github.com/zkcpku/CodeAgent
[…]
CodeAgent: Enhancing Code Generation with Tool-Integrated Agent Systems for Real-World Repo-level Coding Challenges
Kechi Zhang [^1] The two authors share equal contribution.,  Jia Li [^1] The two authors share equal contribution.,  Ge Li [^2] Corresponding authors.,  Xianjie Shi,  Zhi Jin [^2] Corresponding authors. Key Lab of High Confidence Software Technology (PKU), Ministry of Education School of Computer Science, Peking University, China {zhangkechi,lijiaa,lige}@pku.edu.cn, 2100013180@stu.pku.edu.cn, zhijin@pku.edu.cn
1 Introduction
Code generation automatically generates programs for the natural language (NL) requirement. Recent years have seen a trend in tackling code generation tasks with large language models (LLMs), such as CodeLlama Rozière et al. (2023), StarCoder Li et al. (2023), and DeepSeekCoder DeepSeek (2023). To evaluate the effectiveness of LLMs on code generation, many efforts have been performed Zhang et al. (2023b); Luo et al. (2023); Zheng et al. (2023). They demonstrate LLMs have impressive code generation abilities, e.g., CodeLlama Rozière et al. (2023) achieves 55% pass rate on the MBPP benchmark Austin et al. (2021). Despite achieving satisfactory performances, they mainly focus on simple generation scenarios including statement-level and function-level code generation. Statement-level code generation Iyer et al. (2018); Athiwaratkun et al. (2022) aims to output statement-specific source codes. Function-level code generation Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021) predicts independent code that only invokes built-in functions and APIs from third-party libraries. For both scenarios, on the one hand, the length of the generated code is rather short. For instance, the average line of each target code is 11.5 on the most popular benchmark HumanEval Chen et al. (2021). On the other hand, they only generate standalone code units. However, more than 70% functions in the open-source projects are non-standalone Yu et al. (2023). Developers typically write programs based on specific code environments that usually have intricate contextual dependencies.
To address the gap in understanding LLMs’ performance in practical scenarios, we attempt to investigate LLMs in a more pragmatic scenario, i.e., repo-level code generation. We manually construct CodeAgentBench, a benchmark specifically for repo-level code generation, comprising five diverse Python code repositories with a total of 101 functions sourced from Github  [^2] https://github.com/. It provides rich information about the repository, such as documentation and contextual dependency, to help LLMs better understand it. Besides, each repository has a self-contained test suite to efficiently verify the correctness of generated codes. Given a requirement, LLMs generate desired programs that not only meet the requirement but also ensure seamless integration with the current project repository. When tested with CodeAgentBench, we observe that even the most powerful model, GPT-4 GPT-4 (2023), only achieves 21.8% pass rate, indicating a decline in performance compared to standalone code generation. These models struggle with handling complex code tasks, often leading to inaccurate code outputs. Further details and examples are provided in Section 6.2 and Figure 5.
To alleviate this problem, we follow the research line of LLM-based agents Palo et al. (2023); Wang et al. (2023a); Xi et al. (2023) and introduce a novel LLM-based agent framework CodeAgent that leverages external tools to help LLMs in repo-level code generation. With five programming tools, CodeAgent is capable of interacting with the software artifacts, including retrieving useful information, finding existing code symbols in the repository, and handling essential code testing. To guide LLMs to efficiently use tools, we draw on four agent strategies including ReAct, Tool-Planning, OpenAI function, and Rule-based form. Based on agent strategies, LLMs can automatically select suitable tools for each repo-level task, finally provide a comprehensive response. To our knowledge, we are the first to adopt and optimize a agent-based method for complex repo-level code generation tasks.
We conduct extensive experiments for evaluation. We apply CodeAgent to nine powerful open-source and closed-source LLMs with parameter sizes ranging from 13B to 175B to show the universality. Compared to directly generating from LLMs, experimental results on CodeAgentBench reveal that CodeAgent achieves performance improvements ranging from 18.1% to an extraordinary 250% across various LLMs. Further evaluations on well-known function-level benchmark HumanEval Chen et al. (2021) confirm CodeAgent’s versatility and effectiveness in diverse code generation tasks. Remarkably, when compared with commercial products like Github Copilot, CodeAgent stands out, demonstrating superior accuracy. These findings highlight the robust practical capabilities of CodeAgent in code generation community, underscoring its potential to evolve real-world repo-level coding challenges.
In summary, we make the following main contributions:
  • We make an attempt to investigate repo-level code generation, which has crucial worth for understanding LLMs’ performance in practical scenarios.
  • We construct CodeAgentBench, a repo-level code generation benchmark. It has high-quality code repositories and diverse topics.
  • We propose CodeAgent that leverages tools to tackle repo-level code generation. We explore five tools to help LLMs complete the whole generation process and draw on four agent strategies to automatically select suitable tools and generation formats.
  • Experimental results on nine LLMs demonstrate our CodeAgent’s versatility and effectiveness in diverse code generation tasks, highlighting its potential for real-world repository-level coding challenges.
2 Background
2.1 LLMs and Agents for Code Generation
Code generation can automatically generate source codes given an NL requirement, which attracts more and more attention in industry and academia. Nowadays, LLMs have shown impressive capabilities in code generation since they have billions of parameters trained on a large amount of corpus with different training objectives. Early, Lu et al. (2021) adapt GPT-2 Radford et al. (2019) model on the source code, resulting in CodeGPT. Chen et al. (2021) fine-tune GPT-3 Brown et al. (2020) models on code corpus to produce CodeX Chen et al. (2021) and GitHub Copilot Dakhel et al. (2023). Recently, OpenAI  [^3] https://openai.com/ proposes GPT-3.5 and GPT-4 series models (e.g., ChatGPT Chat (2022)), which have shown strong generation capabilities in natural language and programming languages. Since neither CodeX Chen et al. (2021) nor GPT-3.5 GPT-3.5 (2023) is open-sourced, some researchers attempt to replicate them, bringing in CodeGen Nijkamp et al. (2022), StarCoder Li et al. (2023), CodeLLaMA Rozière et al. (2023), WizardCoder Luo et al. (2023) and DeepSeekCoder DeepSeek (2023). Recent research has increasingly shown that LLMs can be instrumental in developing AI agents Palo et al. (2023); Wang et al. (2023a); Xi et al. (2023). Examples such as ToolFormer Schick et al. (2023), HuggingGPT Shen et al. (2023), Gorilla Patil et al. (2023), ToolLlama Qin et al. (2023) and ToolCoder Zhang et al. (2023a) demonstrate LLMs’ proficiency in tool utilization. Projects like Auto-GPT AutoGPT (2023), BabyAGI BabyAGI (2023), and KwaiAgents Pan et al. (2023) highlight the autonomous capabilities of LLM-powered agents in fulfilling human tasks. However, there is no relavent work targeting the complex coding capabilities of agent systems. In this paper, we select GPT-4 GPT-4 (2023), GPT-3.5 GPT-3.5 (2023) and other powerful LLMs to design coding agent systems for real-world repo-level code generation.
2.2 Benchmarks for Code Generation
The code generation benchmarks contain various coding tasks and commonly verify the correctness of generated code by Pass@k Chen et al. (2021). To date, many code generation benchmarks Yin et al. (2018); Athiwaratkun et al. (2022); Hendrycks et al. (2021) focus on generating independent programs. Based on the granularity of examples, these benchmarks mainly be categorized into statement-level and function-level generation. Statement-level generation only predicts statement-specific source codes. CoNaLA Yin et al. (2018) is a representative statement-level benchmark where the target of each example contains one statement. Function-level generation Hendrycks et al. (2021); Chen et al. (2021) requires LLMs to predict standalone function-oriented programs given an NL requirement. The generated programs only invoke built-in functions or APIs from third-party libraries. This is by far the most common code generation task. Besides, some benchmarks Li et al. (2022) contain algorithm questions. LLMs aim to generate an algorithm solution like a human programmer. These benchmarks play a significant role in the advancement of the code generation field.
However, in software development, programmers mainly work within a specific code environment, commonly referring to a code repository. The code repository includes documentation, predefined programs, historical issues, and submitted commits, serving as the foundational framework for developers to build and extend their functionalities. In this case, developers must understand and navigate the code repository environment thoroughly to ensure seamless integration of newly added code with existing code. Recently, some project-level Yu et al. (2023) and repository-level Liao et al. (2023) benchmarks have been proposed to investigate the challenging generation scenario. Nevertheless, they only provide limited constraint information to LLMs, such as the requirements, signature information, and restricted code dependencies. When it comes to LLM-based code generation techniques, solely relying on inputting simplified information to LLMs fails to align with the existing environment. In this paper, we construct a repository-level code generation benchmark CodeAgentBench. Compared with existing repository-level benchmarks, CodeAgentBench provides rich input information (i.e., requirement, documentation, code dependency, and runtime environment), and self-contained test suites. It is closer to real-world programming scenarios, fostering the evolvement of code generation community.
Benchmark Language Source Task Samples # Tests # Line # Tokens # Input
CoNaLA Yin et al. (2018) Python Stack Overflow Statement-level 500 1 4.6 NL
Concode  Iyer et al. (2018) Java Github Function-level 2000 - 26.3 NL
APPS Hendrycks et al. (2021) Python Contest Sites Competitive 5000 21.4 58 NL + IO
HumanEval Chen et al. (2021) Python Manual Function-level 164 11.5 24.4 NL + SIG + IO
MBXP Athiwaratkun et al. (2022) Multilingual Manual Function-level 974 6.8 24.2 NL
InterCode Yang et al. (2023) SQL, Bash Manual Function-level 200, 1034 - - NL + ENV
CodeContests Li et al. (2022) Python, C++ Contest Sites Competitive 165 59.8 184.8 NL + IO
ClassEval Du et al. (2023) Python Manual Class-level 100 45.7 123.7 NL + CLA
CoderEval Yu et al. (2023) Python, Java Github Project-level 230 30.0 108.2 NL + SIG
RepoEval Liao et al. (2023) Python Github Repository-level 383 - - NL + SIG
CodeAgentBench Python Github Repository-level 101 57.0 477.6 Software Artifacts (NL + DOC + DEP + ENV)
Table 1: The statistics of existing widely-used code generation benchmarks. # Tests: whether a benchmark has the test suite. # Line: average lines of code. # Tokens: average number of tokens. # Input: Input information of LLMs. NL: Natural language requirement. IO: Input and output pairs. SIG: Function signature. CLA: Class skeleton as described in Section $2.$ ENV: Runtime environment. DOC: Code documentation. DEP: Code dependency.
3 CodeAgentBench Benchmark
In this section, we introduce our CodeAgentBench benchmark. We illustrate its composition format (Section 3.1) and construction process (Section 3.2) in detail.
Name Domain Samples # Line # DEP
numpyml-easy Machine Learning 22 10.9 0.3
numpyml-hard Machine Learning 35 85.4 2.6
container Data Structure 4 130.3 8.0
micawber Information Extraction 7 19.7 4.3
tinydb Database 21 36.7 2.7
websockets Networking 12 91.6 7.5
Total 101 57.0 3.1
Table 2: Statistics of CodeAgentBench. # Line: average lines of code. # DEP: average number of code dependencies.
3.1 Benchmark Composition
Code repository contains intricate invocation relationships. Only with a deep understanding of code repository can LLMs generate satisfying programs that not only adhere to requirements but also seamlessly integrate with the current repository. Inspired by this, each task of our benchmark provides rich information, encompassing the documentation, code dependency, runtime environment, self-contained test suite, and canonical solution, which form the input and output.
3.1.1 Benchmark Input
Documentation
Documentations are the main input component of our benchmark and describe the generation targets. The documentation provides additional supporting information beyond the NL requirements. It contains target class-level (class name, signature, and member function) and function-level (functional description, and params description) information. Typically, the correctness of generated programs is verified with the test suite. The generated programs must conform to the interface (e.g., the input parameters). Thus, the documentation also provides the type and interpretation of input parameters and output values. Considering that requirements usually contain domain-specific terminologies, the documentation also explains these terms such as the mathematical theorem. We follow the code documentation format used in a popular documentation creation tool Sphinx  [^4] https://www.sphinx-doc.org/. Figure 1 illustrates an example of documentation in CodeAgentBench, where different elements are highlighted with diverse colors. When accomplishing a new task, our prepared documentation can provide LLMs with all-sided details that need to be considered to ensure that the generation target has been well-defined and constrained.
Figure 1: An illustrative example of CodeAgentBench. Our proposed CodeAgentBench contains complex descriptions and code dependencies, which is more realistic than the existing benchmark.
Figure 2: An illustrative example of HumanEval.
Contextual Dependency
A key distinction of our CodeAgentBench from other benchmarks is its inclusion of contextual dependencies. This aspect is crucial, as classes or functions typically interact with other code segments within the same repository, such as import statements or other user-defined classes and functions. These interactions may occur within the same file or across multiple files. For instance, to implement the RandomForest class in Figure 1, it is necessary to utilize the bootstrap_sample function from rf.py and the DecisionTree class from dt.py, demonstrating the intricate code contextual dependencies involved.
To accurately identify these dependencies, we developed a static analysis tool using tree-sitter  [^5] https://tree-sitter.github.io/tree-sitter/. Our designed tool allows us to extract all user-defined elements (such as class names, function names, constants, and global variables) and public library names from each file. These elements are then stored in a knowledge base. For any given function, we use this knowledge base to locate its source file, parse the file to identify all user-defined symbols and public libraries, and finally determine its contextual dependencies by exact matching of symbol names and scopes. On average, each sample in CodeAgentBench involves around 3.1 code dependencies, thereby closely simulating real-world programming conditions. Detailed information is shown in Table 2.
Runtime Environment
Different from natural language, program language is executable. Whether programs return target results after execution is a crucial manner to verify the correctness of generated programs. Developers typically depend on the execution feedback to correct errors in programs. In CodeAgentBench, we build a sandbox environment for each task. The sandbox environment provides all configurations needed to run the repository and offers convenient interaction to ensure an all-sided evaluation of LLMs’ performance on repo-level code generation.
3.1.2 Benchmark Ground-truth Output
Canonical Solution
We use the answers included in the repository as the initial solutions and invite three participants to manually refine them. The first participant checks surface errors of solutions based on the repository information. The second person runs the solutions to identify and fix execution bugs. The last participant is responsible for executing solutions with the test suite, aiming to ensure its functional correctness. Through the iterative process, we can ensure the robustness and reliability of solutions as much as possible.
3.1.3 Benchmark Evaluation
Self-Contained Test Suite
To evaluate the correctness of generated programs, CodeAgentBench furnishes a self-contained test suite for each task. We first analyze and extract test cases contained in the repository. We then invite two participants to manually add test cases to enhance its coverage as much as possible. In CodeAgentBench, each task has at least one unit test case. Whereafter, another participant manually checks the correctness of the test suite. Given a new task, we run the corresponding unit test code to verify the generated programs based on our sandbox environment. We treat the generated program correctly only if its output aligns with all ground truths of the test suite. For fairness, LLMs can not access the test suite during code generation.
3.2 Benchmark Construction Process
To make CodeAgentBench diverse, we select five prevalent topics judged by ten developers and choose repositories with high stars from GitHub. The selected topics contain machine learning, data structure, information extraction, database, and networking. To ensure the quality, we only select repositories that use pytest  [^6] https://docs.pytest.org/ and unittest [^7] https://docs.python.org/3/library/unittest.html as the test framework and its documentation is generated by Sphinx [^8] https://www.sphinx-doc.org/ tool. We also filter out complex repositories that are hard to deploy and test. Then, we extract all functions and classes in repositories and arrange two participants to sequentially execute them. Our construction costs approximately 600 person-hours. Each participant possesses 2-5 years of Python programming experience. Finally, we get 101 functions with 5 repositories in Python. The statistics of CodeAgentBench are shown in Table 2. For the potential data leakage issue, we conduct a preliminary experiment and discussion in Seciton 6.1.
Compared to the previous benchmarks, our CodeAgentBench has two main advantages. On the one hand, it is closer to real-world code generation scenarios. On the other hand, CodeAgentBench provides self-contained information, including documentation, contextual dependency, runtime environments, and test suites. The information can efficiently prompt LLMs for repo-level code generation.
4 CodeAgent Method
We perform a preliminary exploration to assess and improve LLMs’ ability on repo-level code generation. Considering that developers often use various tools to assist with programming, we follow the research line of LLM-based agents Palo et al. (2023); Wang et al. (2023a); Xi et al. (2023) and introduce a novel LLM-based agent framework CodeAgent that leverages external tools to enhance the problem-solving abilities of LLMs in intricate repo-level code generation. CodeAgent seamlessly pauses generation whenever tools are called and resumes generation by integrating their outputs. The tools of CodeAgent can assist LLMs with the entire code generation process, including information gathering, code implementation, and code testing (Section 4.1), thus interact with the software artifacts. Providing LLMs with access to tools, CodeAgent explores four agent strategies to guide LLMs to interact with these tools properly for generation (Section 4.2). Figure 3 illustrates the overview of our CodeAgent.
Figure 3: Overview of our CodeAgent.
4.1 Tools
Given a requirement, developers usually first gather relevant knowledge, then find and modify existing programs to meet the requirement, and finally verify generated programs with the assistance of tools. To mimic this process, CodeAgent incorporates external tools from three perspectives: information retrieval, code implementation, and code testing, which are commonly used by programmers in their daily work.
Tool Domain Tool Name Format
Information Retrieval Website Search WebSearch(input_query)
Documentation Reading DocSearch(input_name)
Code Implementation Code Symbol Navigation SymbolSearch(module_path or input_name)
Code Testing Format Checker FormatCheck()
Code Interpreter PythonREPL(input_code)
Table 3: Tool statistics in CodeAgent
4.1.1 Information Retrieval Tools
Information retrieval tools are responsible for analyzing repositories and collecting resources, which is pivotal in understanding the problem domain. In this paper, we use popular website search and documentation reading tools to gather information.
Website Search
Programmers often share solutions for various programming problems on websites where search engines consider them as knowledge resources. When encountering similar problems, developers only submit a question query to a search engine. The engine can provide useful programming suggestions. Inspired by this, CodeAgent uses a popular search engine DuckDuckGo [^9] https://duckduckgo.com/ to choose the most relevant websites, and then apply LLMs to summarize the website content as the final tool output.  [^10] We choose DuckDuckGo because it provides a cheaper and more convenient API than other search engines such as Google and Bing. In the process, we block websites that may lead to data leakage. This tool is designed in the format: WebSearch(input_query), and the tool will return the formatted content searched from websites. If the result is too long, the tool will use the LLM to summarize it.
Documentation Reading
Besides gathering information from websites, we also retrieve relevant knowledge from the documentation of the repository. To achieve this, CodeAgent leverages BM25 Robertson et al. (2009) as the documentation reading tool. Given a class name or function name, it can retrieve correlative content from the documentation as its output and then provide them to LLMs for code generation. This tool is designed in the format: DocSearch(input_name), and the tool will return the content searched from project documentations. Similar to the website search tool, the tool will summarize the results as necessary.
4.1.2 Code Implementation Tools
Code implementation tools aim to provide relevant code items (i.e., pre-defined symbol names and code snippets) in the repository. LLMs modify and integrate these items into their generation process. It not only expedites the development process but also encourages code reuse. We leverage code symbol navigation as the code implementation tool.
Code Symbol Navigation
Based on tree-sitter  [^11] https://tree-sitter.github.io/tree-sitter/, we build a code symbol navigation tool for Python. This tool explores code items from two types. The first type is file or module-oriented parsing, where code symbol navigation performs static analysis of a file or module and provides symbol names defined in it, encompassing global variables, function names, and class names. The other type is the class or function symbol navigation. Given a class or function name, the tool finds its definition from the code repository. Combining the two types, this tool can traverse predefined source code within a repository, empowering LLMs to understand intricate dependencies and reuse codes. This tool is designed in the format: SymbolSearch(module_path or input_name). The tool will detect what the input is and return the corresponding results (e.g., all defined symbols in the given file path or the implementation code corresponding to the given symbol name). When no parameters are provided, the default value is the path of the current file.
4.1.3 Code Testing Tools
After acquiring generated codes, we use code testing tools to format and test them, enhancing their correctness and readability.
Format Checker
It is employed to check the format correctness of generated codes. Specifically, we use Black  [^12] https://github.com/psf/black as the format checker, which is a Python code formatter. It executes generated codes and checks format errors such as indentation misalignment and missing keywords. Subsequently, it tries to rectify these errors and reorganizes code statements, enhancing the correctness and readability of generated codes. This tool is designed in the format: FormatCheck(), and the tool will automatically formats the model’s most recently generated code and return the formatted version.
Code Interpreter
The tool focuses on examining the syntax and function of programs. It furnishes a runtime environment so that LLMs can debug generated codes with execution feedback. The tool requires LLMs to provide a program to be executed, and then runs the code in the repository environment. Meanwhile, LLMs generate some test cases to verify whether the output of the generated code meets the expected results. When occurring errors, this tool will offer error information to facilitate LLMs to fix bugs until programs are error-free, which has been proven to be effective by many existing works Chen et al. (2022); Zhang et al. (2023b) to improve the accuracy of output programs. The runtime environment is prepared for each task, as described in Section 3.1.1. This tool is designed in the format: PythonREPL(input_code), and the tool will return the executed result of the input code.
4.2 Agent Strategy
To guide LLMs to leverage these powerful tools properly, we explore three existing agent strategies for repo-level code generation, including ReAct Yao et al. (2022), Tool-Planning Wang et al. (2023b), and OpenAIFunc OpenAI-Function (2023). These agent strategies are put forward to augment LLMs’ reasoning and problem-solving abilities. In addition, inspired by the human programming process, we also propose a Rule-based strategy to access with tools. We perform our experiments based on LangChain  [^13] https://python.langchain.com.
ReAct
The strategy is a general paradigm that synergizes reasoning and acting with LLMs, while also interacts with the external environments to incorporate additional information for solving problems Yao et al. (2022). It allows LLMs to perform dynamic reasoning and thus can adjust and maintain reasoning plans in time. In repo-level code generation, ReAct prompts LLMs to generate reasoning traces and task-related actions in an interlaced fashion. Based on actions, ReAct selects the proper external tools and invokes them by providing input. The strategy then treats the output of tools as additional knowledge and decides whether to generate a final code or invoke other tools for further processing.
Tool-Planning
We propose a variant, i.e., Tool-Planning, of Planning strategy Wang et al. (2023b) that makes a plan before solving problems and has shown effectiveness in many studies Zhang et al. (2022); Jiang et al. (2023). Different from Planning, our strategy can invoke proper tools based on the plan. Specifically, Tool-Planning first makes a plan to divide an entire task into several subtasks and then performs subtasks according to the plan. For complex subtasks, it will automatically choose appropriate tools to assist LLMs in code generation.
OpenAIFunc
Recently, some models (e.g., GPT-3.5 GPT-3.5 (2023) and GPT-4 GPT-4 (2023)) have the function-calling ability. Given a task description, they can detect when a function should be called, thus rendering them ideal for integration with external tools. In this strategy, we use the function-calling interface OpenAI-Function (2023) provided by OpenAI, which can accept human-created tools. For repo-level code generation, we provide the model with external tools described in Section 4.1. It then determines the time and type of invoking tools, and offers the input to the called tools. LLMs combine with the returned results of tools to generate programs.
Rule-based Tool Usage
When faced with a complex problem, programmers often first learn external knowledge (e.g., website search) and background knowledge within the repository, then write programs, and finally check the functionality of programs. Inspired by the workflow, we propose a rule-based strategy for generating codes.
The strategy defines the order of tool usage and interlinks these tools through prompts. ❶ LLMs leverage website search to gather useful online information; ❷ LLMs then use documentation reading tool to search relevant classes and functions; ❸ code symbol navigation is required to select and view the source code of related classes and functions, meanwhile learn the pre-defined symbols in the same file as the target code. Based on the above information, LLMs generate programs. ❹ Subsequentially, LLMs invoke the format checker to check the syntax and format of generated programs. ❺ Finally, LLMs generate some test cases and use the code interpreter to evaluate the functional correctness of programs. LLMs fix the errors until the feedback information is error-free. For each part, the large model will autonomously cycle through the use of tools until it decides to move on to the next part or the cycle reaches its limit (we set the limit to be less than 3).
5 Experiment
Based on the constructed benchmark and agent strategies, we conduct the first study to evaluate existing LLMs on repo-level coding challenges. We perform extensive experiments to answer four research questions: (1) How much can CodeAgent improve the advanced code generation LLMs on CodeAgentBench (e.g., repo-level code generation) (Section 5.2); (2) What is the improvement of our CodeAgent on classical code generation such as HumanEval (Section 5.3); (3) To what extent do our selected tools in the agent system help for repo-level coding (Section 5.4); (4) How helpful is CodeAgent for human programmers compared with the commercial code generation products (Section 5.5).
5.1 Experiment Setup
Dataset
We evaluate our approach on our new benchmark CodeAgentBench as described in Section 3. For each task in the benchmark, LLMs are provided with documentation containing the requirements needed to be implemented, along with a set of tools we designed, as well as full access permissions to code files in the repository. In addition, to evaluate the generalization ability of our CodeAgent, we also perform experiments on function-level code generation. In this paper, we use HumanEval Chen et al. (2021) since it is a widely used benchmark. It contains 164 programming problems with the function signature, docstring, body, and unit tests. We give an illustration of these benchmarks in Figure 1 and Figure 2.
Base LLMs
We investigate the effectiveness of several popular code LLMs. Specifically, we apply CodeAgent to nine most powerful LLMs, including GPT-3-davinci GPT-3 (2022), GPT-3.5-turbo GPT-3.5 (2023), GPT-4-turbo GPT-4 (2023), Claude-2 Claude (2023), Llama2-70B-chat Llama (2023), CodeLlama-34B Rozière et al. (2023), WizardCoder-34B Luo et al. (2023), DeepSeek-33B DeepSeek (2023) and Vicuna-13B Chiang et al. (2023). Additional descriptions are provided as a part of Table 4.
Metrics
Following previous works Zan et al. (2022); Zheng et al. (2023), we use the pass rate as the metric to evaluate the performance of LLMs, where we treat the generated program correctly only if its output is consistent with all ground truths of the test suite. Specifically, we are mainly concerned with Pass@1 Chen et al. (2021), which is a representative of the Pass@k family, because in real-world scenarios, we usually only consider the single generated code.
5.2 Repo-level Coding Task
Models Scales NoAgent Rule-based ReAct Tool-Planning OpenAIFunc
Closed source LLM
GPT-3-davinci GPT-3 (2022) 175B 16.8 24.8 ( $\uparrow$ 7.9) 22.8 ( $\uparrow$ 5.9) 18.8 ( $\uparrow$ 2.1) -
GPT-3.5-turbo GPT-3.5 (2023) - 19.8 31.7 ( $\uparrow$ 11.9) 30.7 ( $\uparrow$ 10.8) 21.8 ( $\uparrow$ 2.0) 28.7 ( $\uparrow$ 8.9)
GPT-4-turbo GPT-4 (2023) - 21.8 37.6 ( $\uparrow$ 15.8) 34.7 ( $\uparrow$ 12.9) 25.7 ( $\uparrow$ 4.0) 34.7 ( $\uparrow$ 12.9)
Claude-2 Claude (2023) - 8.9 10.9 ( $\uparrow$ 2.0) 9.9 ( $\uparrow$ 1.0) 9.9 ( $\uparrow$ 1.0) -
Open source LLM
LLaMA2-70B-chat Llama (2023) 70B 10.9 12.9 ( $\uparrow$ 2.0) 11.9 ( $\uparrow$ 1.1) 11.9 ( $\uparrow$ 1.1) -
CodeLLaMA-34B Rozière et al. (2023) 34B 2.0 5.0 ( $\uparrow$ 3.0) 4.0 ( $\uparrow$ 2.0) 4.0 ( $\uparrow$ 2.0) -
WizardCoder-34B Luo et al. (2023) 34B 2.0 6.9 ( $\uparrow$ 5.0) 5.0 ( $\uparrow$ 2.7) 4.0 ( $\uparrow$ 2.0) -
DeepSeek-33B DeepSeek (2023) 33B 13.9 24.8 ( $\uparrow$ 10.9) 20.8 ( $\uparrow$ 6.9) 15.8 ( $\uparrow$ 2.0) -
Vicuna-13B Chiang et al. (2023) 13B 1.0 1.0 0.0 0.0 -
Table 4: The Pass@1 results of different agent strategies on CodeAgentBench. "NoAgent" refers to the baseline where LLMs generate code solely based on the provided documentation. Other columns indicate our agent strategies as described in Section 4.2.
In our experiments, we utilized CodeAgentBench to assess the efficacy of CodeAgent in enhancing the performance of nine prominent code LLMs. The specially designed benchmark, CodeAgentBench, is repository-level code generation that encompasses intricate documentation and code dependencies, accurately mirroring real-world scenarios. The results of applying CodeAgentBench to these nine models are presented in Table 4.
Our proposed CodeAgentBench proves to be substantially more challenging than existing benchmarks, as evidenced by the relatively lower pass rates. On all base LLMs with various sizes, CodeAgent consistently delivers significant performance improvements. Specifically, for GPT-4 model GPT-4 (2023), we observe a maximum increase of 15.8%, equating to a 72.7% relative enhancement over the baseline, i.e., NoAgent. The improvements of other LLMs range from 18.1% to an impressive 250%, underscoring the effectiveness of our proposed approach. This demonstrates that the tools integrated within CodeAgent provide crucial repo-level information, aiding LLMs in producing accurate code solutions and effectively tackling complex repo-level coding challenges.
Across different LLMs, a notable trend is that more advanced LLMs exhibit greater improvements with the application of CodeAgent. However, for Vicuna-13B model Chiang et al. (2023), performance on CodeAgentBench is notably poor, showing no appreciable enhancement with the agent strategy. In contrast, the improvement is quite pronounced for other high-capacity LLMs. Furthermore, we find that different agent strategies yield varying levels of enhancement. Among these strategies, Rule-based and ReAct strategies are more effective, whereas Tool-Plannig strategy appears less suited for the task.
5.3 Function-level Coding Task
Models NoAgent Rule-based ReAct Plan OpenAIFunc
GPT-3.5-turbo GPT-3.5 (2023) 72.6 82.3 ( $\uparrow$ 9.7) 79.3 ( $\uparrow$ 6.7) 73.8 ( $\uparrow$ 1.2) 81.1 ( $\uparrow$ 8.5)
CodeLLaMA-34B Rozière et al. (2023) 51.8 59.7 ( $\uparrow$ 7.9) 58.2 ( $\uparrow$ 6.4) 54.1 ( $\uparrow$ 2.3) -
WizardCoder-34B Luo et al. (2023) 73.2 79.4 ( $\uparrow$ 6.2) 77.6 ( $\uparrow$ 4.4) 75.6 ( $\uparrow$ 2.4) -
DeepSeek-33B DeepSeek (2023) 78.7 84.8 ( $\uparrow$ 6.1) 83.5 ( $\uparrow$ 4.8) 81.1 ( $\uparrow$ 2.4) -
Table 5: The Pass@1 results of different agent strategies on the HumanEval benchmark.
We further apply our CodeAgent to function-level code generation with the well-known HumanEval benchmark Chen et al. (2021). We adapt our approach to this scenario by omitting the documentation reading tool and code symbol navigation. The adjustment is necessitated as these tools are not applicable to the standalone code generation task. For this task, we strategically selected a range of representative LLMs for evaluation, constrained by our available resources and computational capacity. The pass rate results are detailed in Table 5.
The results once again highlight the efficacy of CodeAgent in enhancing the performance of code LLMs across all metrics. Notably, the maximum improvements observed for each model span from 7.8% to 15.3%. These findings underscore the versatility and effectiveness of our proposed CodeAgent in augmenting the capabilities of LLMs across a variety of code generation tasks.
5.4 Ablation Study
# Usage Ablation Result
GPT-3.5-ReAct - 30.7
Websit Search 0.30 27.7 ( $\downarrow$ 3.0)
Documentation Reading 0.84 26.7 ( $\downarrow$ 4.0)
Code Symbol Navigation 2.45 22.8 ( $\downarrow$ 7.9)
Format Check 0.17 29.7 ( $\downarrow$ 1.0)
Code Interpreter 0.22 29.7 ( $\downarrow$ 1.0)
GPT-3.5-NoAgent - 19.8
Table 6: Average Tool Usage Number and Ablation Result on CodeAgentBench for GPT-3.5-ReAct.
To investigate the influence of tools incorporated in CodeAgent, we conduct an ablation study focusing on tool utilization in repo-level code generation. We choose GPT-3.5-turbo with ReAct as the base model, named GPT-3.5-ReAct. We meticulously track the usage frequency of each tool during code generation processes, with the statistics presented in Table 6 under the column # Usage. Subsequently, we exclude one tool at a time from our approach, allowing us to isolate and understand the individual contribution of each tool. The performances of these ablation scenarios are shown in Table 6, categorized under the column Ablation Result.
Our findings reveal that the code symbol navigation tool is particularly pivotal in our agent system. On average, CodeAgent utilizes this tool approximately 2.45 times per code generation, a frequency higher than the counterpart of other tools. Notably, the performance significantly declines when this tool is omitted, underscoring its critical role in enhancing the effectiveness of our approach. Furthermore, the ablation results confirm that each tool in our agent system contributes positively to the overall improvement. This evidence not only validates the effectiveness of our strategy design but also highlights the utility of programming tools in addressing the repo-level coding task.
5.5 Compared with Commercial Products
NumpyML-easy NumpyML-hard
Our Agent
GPT-3.5 14 3
GPT-4 17 5
IDE Product
GitHub Copilot 7 1
Amazon CodeWhisperer 5 0
Agent Product
AutoGPT (with GPT-4) 2 0
Table 7: Performance compared with commercial programming products (the number of solved problems).
Nowadays, a lot of mature commercial products are available to support complex code generation tasks. It is essential to compare CodeAgent with these established products. We categorize them into two distinct groups: (1) IDE Products are AI-powered autocomplete-style suggestion tools integrated within IDE software. Notable examples are Github Copilot [^14] https://github.com/features/copilot and Amazon CodeWhisperer [^15] https://aws.amazon.com/codewhisperer/. (2) Agent Products encompass autonomous agents driven by GPT-4 GPT-4 (2023). They are capable of executing a variety of tasks, including coding, such as well-known AutoGPT AutoGPT (2023).
We employ the versions of these products with their default settings. Considering that IDE products are primarily designed as completion systems, we limit human interactions to less than three times per task to ensure a fair comparison. The evaluation is conducted on the numpyml subset of CodeAgentBench. All tasks are manually executed by an experienced Python developer. Table 7 shows the number of solved problems on different products and our CodeAgent.
The results demonstrate that CodeAgent works better than existing products on complex coding scenarios. In addition, despite both CodeAgent and AutoGPT being agent-based approaches, CodeAgent exhibits numerous optimizations tailored for repo-level coding tasks, thereby making it better than AutoGPT in the task. Compared to IDE products that can also analyze complex code dependencies, our approach benefits from the flexibility inherent in the agent system, which can use different tools whenever needed, resulting in a substantial lead over IDE products. These findings show the practical capabilities of CodeAgent in the code generation community.
6 Discussion
6.1 Impact of LLMs’ Memorization on Pre-Training Data
To examine the potential impact of LLMs memorization on pre-training data, we conduct a simplified obfuscation study by extending the methodology Tang et al. (2023). Specifically, we modify the input documentation with obfuscated prompts. These prompts are altered by replacing function or parameter names with placeholder terms, which are concise GPT-generated summaries in their original form. We also reorganize the file structure of the code repository. The adjustment aims to challenge LLMs’ memorization on pre-training data.
The results of the obfuscation study, while slightly lower, are comparable to the counterpart obtained with non-obfuscated versions. The obfuscated prompts do not hinder LLMs’ code generation ability. The findings suggest that the performance of LLMs on repo-level code generation cannot be solely attributed to memorization. On the contrary, our propose CodeAgent is the key to performance improvement. We plan to conduct additional experiments on our CodeAgentBench to further investigate the extent of LLMs’ memorization on pre-training data.
6.2 Case Study
We further observe the generated cases to assess CodeAgent (e.g., GPT-3.5-ReAct) and the baseline model (e.g., GPT-3.5-NoAgent). The comparative analysis is shown in Figure 4 and Figure 5.
We can find a distinct operational pattern in CodeAgent. It typically begins with examining the code dependencies in the repository, subsequently refining its code generation strategy through a logical, step-by-step process known as “chain-of-thought”. As in Figure 4, the input documentation specifies the need for a class with member functions set_params and summary. CodeAgent, assisting with the symbol navigation tool, finds the base class and identifies the member function _kernel as a key component for implementation. This is reflected in the generated thought process:
"The set_params and summary methods can be inherited from the base class without modifications … The ‘_kernel’ method needs to be overridden …"
(Generated by CodeAgent-GPT-3.5-ReAct)
On the contrary, GPT-3.5-NoAgent lacks access to detailed information on code structures, resulting in incorrect code solutions, as depicted in Figure 5.
Figure 4: Case Study of our proposed CodeAgent (GPT-3.5-ReAct) on CodeAgentBench. We omit some input prompts and LLM’s thoughts for viewing convenience.
Figure 5: Case Study of the baseline GPT-3.5-NoAgent on CodeAgentBench. Compared with Figure 4, GPT-3.5-NoAgent generates the incorrect programs.
7 Conclusion
We explore how LLMs perform in the real-world repo-level code generation task. We create CodeAgentBench, a new benchmark for repo-level code generation that includes rich information about the code repository, such as documentation and contextual dependency, to help LLMs better understand the code repository. To enhance the repo-level code generation ability, we propose CodeAgent, an agent-based framework that uses external tools to assist with LLMs. Evaluation on nine LLMs shows that CodeAgent works well for diverse programming tasks, highlighting its potential in real-world repo-level coding challenges.
Limitation
Although our work is a very early exploration of this area, there are several limitations on our work that we aim to address as quickly as possible:
Firstly, we propose a new task format for the repo-level code generation task and release CodeAgentBench. Our preliminary experiments prove that the impact of LLMs’ memorization on pre-training data is slight for fair evaluation. However, it still needs further experiments to eliminate this hidden danger. We will follow the relevant research to further understand its influence on our proposed benchmark.
Secondly, we only incorporate simple tools to CodeAgent. Some advanced programming tools are not explored. The limitation may restrict the agent’s ability in some challenging scenarios.
Thirdly, in Section 5.5, the comparison with commercial products is not rigorous since experiments are done manually. We will study how to evaluate IDE products more standardly.
Finally, since LLMs are very sensitive to input prompts, it is very important to optimize prompts in the agent system. We will continue to explore better agent strategies based on the current approach.